### v1

# Whether to trust: the ML leap of faith

Abstract
Human trust is critical for trustworthy AI adoption. Trust is commonly understood as an attitude, but we cannot accurately measure this, nor manage it. We conflate trust in the overall system, ML, and ML’s component parts; so most users do not understand the leap of faith they take when they trust ML. Current efforts to build trust explain ML’s process, which can be hard for non-ML experts to comprehend because it is complex, and explanations are unrelated to their own (unarticulated) mental models. We propose an innovative way of directly building intrinsic trust in ML, by discerning and measuring the Leap of Faith (LoF)taken when a user trusts ML. Our LoF matrix identifies where an ML model aligns to a user’s own mental model. This match is rigorously yet practically identified by feeding the user’s data and objective function both into an ML model and an expert-validated rules-based AI model, a verified point of reference that can be tested a priori against a user’s own mental model. The LoF matrix visually contrasts the models’ outputs, so the remaining ML-reasoning leap of faith can be discerned. Our proposed trust metrics measure for the first time whether users demonstrate trust through their actions, and we link deserved trust to outcomes. Our contribution is significant because it enables empirical assessment and management of ML trust drivers, to support trustworthy ML adoption. Our approach is illustrated with a long-term high-stakes field study: a 3-month pilot of a sleep-improvement system with embedded AI .
1 Introduction
Lack of trust is often cited as a barrier to artificial intelligence (AI )adoption, especially in machine learning (ML). Trust has always been a critical enabler of new technology adoption: people tend to rely on automation they trust, and shun automation they distrust, in the real world [1] and the laboratory [2]. However, ML faces some unique challenges, which are discussed in the context of the psychology, management, and computer-science literatures.
Trust is commonly understood as a unitary concept but, in practice, it is complex. Lee and Sees’ human-centred definition is most commonly accepted: the attitude that the technology will help us achieve our goals [3] in risky circumstances [4], i.e. in an uncertain and vulnerable situation [5]. Initial trust will vary for a technology because we each have a different general tendency to trust due to factors like our culture or personality (‘dispositional’ trust); and a different ability to deal with the situation and the technology (‘situational’ trust) [6]. Once we experience the specific technology, we can develop ‘learned’ trust. Most of the discussion of trust in the literature relates to cognitive or rational drivers of which the user is aware; but trust can also be emotional [7] and/ or have automatic or unconscious drivers, like impulses from learned associations and innate biases [8]. Lee and Sees’ ‘performance’/ ‘process’/ ‘purpose’framework [3] encapsulates most conscious drivers, but misses automatic ones and does not illuminate emotional ones. They argue potential for trust in technology is highest when: the trustee perceives ‘performance’ as strong; the trustee understands how the ‘process’ operates [9]; and the trustee believes its ‘purpose’ is aligned to their own goals. This framework is now used to reflect on the trust challenges faced by ML and tools to manage trust.
ML should, in theory, garner higher trust when performance is better than alternatives. For example, a deep-learning model outperformed a team of 6 radiologists by a statistically-significant 11.5% in a 27,367-woman UK and US breast-cancer study, measured by relative area under the ROC curve [10]. However, if these technologists tested model output against what they expected given the inputs, they would disagree on outputs where the model outperformed them: false positives and negatives were 3.9-15.1% (UK-US) higher than the model. Despite this ML model’s performance, 4 years later, the NHS in the UK does not yet routinely use ML for breast-cancer screening, with trials still ongoing.
The Explainable AI (XAI)movement has been motivated by a desire to improve our understanding of the ‘process’ of how ML models’ algorithms work [11]. However, because ML works by finding the best fit for mathematical models with large data sets [12], it can be hard to follow its reasoning process. Tools like feature importance or Shapley causal quasi values [13] reflect the factors that contribute to an ML model’s conclusion a posteriori [14], not its reasoning. If we understand a reasoning process and it matches our prior knowledge of sensible reasoning, we can directly build intrinsic trust; if not, only extrinsic trust is possible: we need external reassurance from an expert or an evaluation [4]. XAI’s efforts will only increase intrinsic trust if experts can explicitly compare the ML model’s reasoning to their own mental model [4]. This is problematic because current explanations do not always fit recipient mental models (their internal representation of how the world works [15]). For example, clinicians can think in terms of evidence-based mechanism of action [16]. Moreover, each human-in-the-loop may have a unique a priori mental model, which we only start to discover when ML outputs do not align with their implicit expectations. When time to resolve discrepancies is limited, intrinsic trust will not be established. Human systems of trust in high-stakes environments can rely on a trusted senior colleague pressure-testing a recommendation. Similarly, an expert human-in-the-loop can test whether a rules-based AI model reflects their mental model, and build intrinsic trust: they can interrogate the logic, testing whether it has been correctly implemented. When ML outperforms humans, especially on large data sets that are cognitively challenging to process, there should always be a mismatch between the model’s reasoning and expert priors, because new relationships are identified. It is thus not possible to directly build intrinsic trust in ML with experts through a collaborative process [17]. Trust can, at best, be extrinsic, relying on external assurances, so even experts require a leap of faith to act: from what they expect to what the model outputs. This leap can be challenging for experts with high self-confidence in their own abilities: they can ignore ML advice even though they know doing so lowers their performance [18].
However, understanding an ML model’s process requires comprehension of more than the algorithm: data and objective function are also critical. These could be understandable, which may require visualisation and education. An interpretable text feature space is: phrased in everyday language avoiding codes (‘readable’); worded so they are quick and simple to absorb (‘human-worded’); reliant on real-world concepts (‘understandable’) [19]. However, this approach does not include target variables, objective function, nor data visualisation, which can aid user when data sets are large.
If a technology’s purpose is clearly communicated, it is generally trusted to do what it was built to do [3], although there could be concerns about how well it might do so. Because humans do not directly control how an ML algorithm reaches a conclusion, ML represents a fundamental shift in agency between humans and technology [20]; especially if outcomes are automated. If intent is not communicated, lack of agency could exacerbate concerns about purpose alignment, as could more general and/ or multi-agent ML: it will be harder to disentangle how its purpose relates to an individual’s own goals, or to trust it will make appropriate trade-offs, given other priorities.
Most of the literature on AI trust-management tools relates to rational drivers, namely performance, process and purpose. Glikson et al. explore tools with more emotional drivers: embodiment and immediacy behaviours [21]. Jacovi et al. suggest if we base our trust on such factors that are unrelated to model trustworthiness, our trust is undeserved [4]. However, a trustworthy system could ethically employ such techniques, e.g. if unconscious drivers preclude acceptance. It is harder to tell whether a system contains AI if the AI is embedded, or part of a larger system: the role of AI is less tangible and therefore harder to perceive, even if its presence is disclosed. Humans find embodied AI  – e.g. physical robots [22] – more tangible and thus easier to trust [21]. Trust in robots starts higher for the same level of model intelligence and deteriorates over time as performance is evaluated; whereas for embedded AI trust starts lower and builds over time [21]. However, underlying this insight is an issue with most trust measurement, especially embedded ML: it is trust in the system that is usually measured, not trust in the ML algorithm as distinct from data or objective function. Immediacy behaviours, like personalisation, which increase perceptions of interpersonal closeness [23], can improve perceived performance and purpose alignment independent of model trustworthiness. For example, psychological tailoring of advertising based on openness-to-experience or extroversion – assessed through digital footprint (e.g. Facebook likes) – increased 3.5 million users’ clicks by 40% and purchases by 50% [24]. Because AI is digital, it is easy to deploy users’ data for this purpose, and generative AI enables tailoring of tone. However, immediacy behaviours can backfire, especially when AI is embedded, reducing perceived decisional freedom and purpose alignment. Uber driver ‘algorithmic management’, with constant individual performance evaluation and feedback, increased driver awareness of the extent of ongoing monitoring, making them resentful of being micro-managed by an algorithm whose purpose was perceived not as not aligned to their goals [25].
We lack a hierarchy of trust drivers and tools to support trustworthy ML adoption. A critical issue is our inability to directly measure attitudinal trust, so we cannot objectively assess what supports it. We can only measure declared, demonstrated and, with our proposed metric, deserved trust, which we now discuss; see Figure 1 for a diagram of how these fit together.
Figure 1: Declared, demonstrated and deserved trust, and how they differ from attitudinal trust. All human-centred concepts are shown in yellow; technology ones in blue; combined in green. Icons by Angriawan Ditya Zulkamain, Arif Hariyanto, Karen Tyler and PEBIAN from thenounproject.com.
Declared trust, what people say they trust, has been the primary focus for measurement [26]. This is derived from attitudinal trust because it only reveals what the respondent is conscious of and is prepared to disclose, given social norms. It will therefore miss automatic drivers of which the respondent is unaware, like biases. Declared trust is analysed using subjective data from interviews and surveys, using different scales so it is hard to learn across studies [27]. Crucial trust drivers may be missing because scales are researcher defined.
Demonstrated trust is a behaviour: we take a risk and act, displaying ‘reliance’ [3]. Demonstrated trust reflects both conscious and automatic drivers but it is not the same as attitudinal trust, e.g. a user may trust but not act because they lack behavioural control, or fear being judged by society; or they might set an intention to act, but not follow-through because of constraints encountered [28]. Demonstrated trust can be objectively measured, e.g. Weight on Advice [18] or percentage of recommendations relied upon [29]. However, these metrics fail to reflect action follow-through, which is necessary when there are constraints to be overcome. Furthermore, there is little exploration of trust-management-tool implications, e.g. there could be a difference between what boosts declared trust and follow-through to action. The relationship between declared trust and action appears strongest when there is high cognitive complexity, e.g. complex automation and novel situations [3]. Reliance can also be stronger when the individual has high decisional freedom as to subsequent action and is able to compare the technology’s recommendation to the human alternative [6].
Warranted or deserved trust is when the risk was worth it because the technology itself was trustworthy [3]. Trustworthiness is a property of technology, not an individual, so it is not necessary for attitudinal or demonstrated trust [3]. Trustworthiness is ideally calibrated to trust so technology is not misused, disused or abused, reducing safety risks [30]. There is substantial discussion on trustworthiness measurement, e.g. the EU criteria include human agency and oversight; transparency; diversity, non-discrimination, fairness; as well as prevention of harm: technical robustness and safety; privacy and data governance; accountability; and societal and environmental well-being [31]. However, the literature on deserved-trust measurement is limited. Jacovi et al. suggest model trustworthiness should be manipulated so it is no longer calibrated with attitudinal trust, to see how much attitudinal trust changes [4]. This is a conceptual approach that is challenging to put into practice in the real world in a verifiable, controlled, reproducible, comparable manner. This approach fails to assess what we really want to know: did trusting the recommendation deliver better outcomes? While a model might be technically robust, it might not take into account factors that would influence an individual’s perception of behavioural control or situation-specific constraints. In anything involving behaviour change, deserved trust is derived from demonstrated trust, not attitudinal, because a recommendation needs to be acted on for impact to be assessed. The individual’s experience of the specific situation and technology, learning whether trust was deserved or not, then translates back to attitudinal trust.
Most users do not therefore understand the leap of faith they take when they trust ML. Current efforts to build trust fail because explanations are unrelated to domain experts’ (unarticulated) mental models. We propose a more practical way of building trust, that shortens the leap. We now illustrate this using data from a long-term high-stakes field study: a 3-month pilot for a sleep-improvement system with embedded AI . We briefly describe our study methods in Section 2, before we propose our novel architecture in Section 3. Section 4 evaluates the contribution of this architecture and its limitations.
2 Methods
The study was high stakes because it impacted health [32], and the scale of changes were substantial. Most participants were short sleepers – i.e. they slept less than 7 hours a night – which is associated with poor health outcomes, e.g. 1.2-1.4x higher Alzheimer’s disease risk [33]. Immune response suffers, so short sleepers are 5x more susceptible to infections like the common cold [34]. Reaction times [35] decline, similar to alcohol intoxication [36]. The main treatment for those who are not chronic insomniacs is ‘sleep hygiene’: up to 16 behavioural and environmental suggestions (e.g. avoid bright light in the evening). While these interventions have a medium effect in healthy-adult sleep trials [37], research has been subjective and failed to take into account individual sensitivity, even though it varies greatly: e.g. one person’s sensitivity to evening light can be 40x that of another person [38]. Slow-Wave Sleep (SWS)declines as we age [39] and clears beta amyloid plaques, linked to Alzheimer’s disease, from the brain [40]. There has been limited research into the impact of these practices on healthy-adult SWS. The situation was thus uncertain and participants vulnerable, so they took a risk when they followed a recommendation.
[…]
Figure 2: Socio-technical sleep-improvement system with embedded parallel neuro-symbolic AI.
Our models were designed to be trustworthy using the EU criteria [31]: human agency and oversight, and transparency are central our socio-technical-system design. We implemented the other principles as follows: a) Models were run on each individual’s own data. b) Participants were informed of model technical robustness and safety: we informed them that rules-based model recommendations were sleep-expert approved and of their own ML model accuracy: 90% (87-94% range, 2.4% SD); and 80% for SWS 80% (75-84% range, 3.3% SD). c) Privacy and data standards were specified by the University Ethics Committee (PREC reference number 22 003, 5th September 2022). These standards were a key selection criteria for smart devices, requiring the team to create their own environmental monitor because no third-party solution was identified that delivered sufficient granularity and met these standards. d) The University Ethics Committee holds the researcher accountable for model trustworthiness and upholding standards, and each participant was accountable for their own data, objective function, and choices. e) The key societal challenge was ensuring participant sleep anxiety did not increase as a consequence of the study. Sleep anxiety was measured before and after the study, and it (statistically insignificantly) declined. Severe anxiety or depression were screened out, through GAD7 and PHQ9 questionnaires. Sleep end points were continually monitored, with an 85% floor on Sleep Efficiency (SE), the percentage of time participants were asleep while in bed.
3 Proposed architecture
Model-architecture functional compartmentalisation enables user oversight over data and objective function, increasing human agency. We use an expert-validated rules-based AI model to create a verified point of reference that can be tested a priori by ML model users. Our LoF matrix visually contrasts this reference standard to our ML model’s output, using the same data and objective function. Areas of agreement can be readily trusted and what is unique to the ML model can be identified and addressed. We propose three simple objective metrics to measure demonstrated trust, distinguishing intention-setting and follow-through, and whether trust was deserved.
3.1 Enabling human agency and oversight
Model-architecture functional compartmentalisation is used to separate data and objective function from model reasoning. Each user was responsible for pre-validating their own data. This exercise uses text and data visualisation techniques, extending the text-driven interpretable-feature-space approach of Zytek et al. [19] to include target variables and the addition of a visual dimension, to make the data more accessible to users who benefit cognitively from visual processing of data.
Validation started with data quality – percentage of each day with complete data for each device – which the system reported daily and weekly against a pre-briefed format and target (70%). Participants progressed to baseline, the next phase of the study, if their one-week trial data was over target.
Participants then assessed whether the data collected during their trial week reflected their perception. Each feature input was provided in a text and visual format that had been reviewed by sleep experts and by 35 participants to ensure each was readable, human-worded and understandable, e.g.: ‘You consume 152mg of caffeine a day (0-286mg range)’, and the chart on the left-hand side of Figure 3.
Figure 3: Caffeine data validation example: upfront data-quality trial chart on the left; baseline chart on the right. The rules embedded in the rules-based model were used to colour-code the latter – in this example, blue for <230mg; red for >400mg; yellow for in-between. Blue was used, rather than green, to make the charts accessible to colour blind participants.
Participants were educated on how to read each chart, what the units meant, and how they were calculated. Participants were encouraged to challenge data accuracy. Some data required considerable interrogation and education because the data and/ or the units were unfamiliar, especially if they were asleep when the data was recorded (e.g. temperature and illumination, as shown in Figure 4).
Figure 4: Unfamiliar-data example recorded while the participant was asleep: bedside temperature on the left; bedside illumination on the right. The nightly average is shown as a horizontal mid-blue line.
Participants trialed different behaviours for unfamiliar items to see how the charts responded. Some changed how an input was recorded if the issue was a diary-input mistake. These follow-ups continued until participants were comfortable their data was accurate.
Participants then reviewed at least 1 month of complete, quality days of baseline data, to confirm target variables and feature inputs were accurate. This included text and charts with sleep-expert-validated logic from the rules-based model, e.g.: ‘You had 330mg of caffeine/ day (blue line). You were over 230mg, the amount which impacts the average person’s sleep, on 82% of days (yellow); over NHS’s 400mg recommended max on 11% of days (red).’ and the chart on the right-hand side of Figure 3.
Each participant had unique sleeping challenges (e.g. waking in the night or too early or having trouble sleeping at night) and objectives: some already achieved well over 7 hours’ sleep and just wanted to improve SWS; some only got 6 hours of sleep and wanted to focus on TST. Sleep objectives were set after the 1-week trial and confirmed once they validated their baseline data and were educated on what was realistic. They thus set model objective function, increasing their agency.
Participants were given the opportunity to remove any baseline data that they did not believe was accurate, but none did. This is an example of demonstrating trust in the data component of an ML model, and of a non-expert being able to build intrinsic trust in part of a complex ML model, because they are able to understand the data inputs and objective function, albeit with support. Data validation, with custom question responses, demonstrated personalisation, an immediacy behaviour, and increased model trustworthiness by ensuring correct inputs.
3.2 Discerning the ML leap of faith
Figure 5: LoF matrix for a pilot participant.
The system recommendation was positioned and visually depicted as an intervention ‘menu’ using a LoF matrix which, as shown in Figure 5, discerns the leap of faith required to trust the ML algorithms. By using a separate a priori expert-validated rules-based AI model to create a reference standard, we contrast the ML models’ prioritisation of which sleep-hygiene changes would make the most difference to the user’s sleep objective, with what an expert would conclude from the same inputs. The user is able to see where the models agree, and where they do not.
Distance from best practice, assessed by the rules-based model, is used to prioritise opportunities on the vertical axis; ML-model feature importance is used to prioritise them on the horizontal. The further away an individual was on best practice, the more impact changing that factor could have on the individual’s sleep. The rules-based model assigned 1 of 3 levels of opportunity to each intervention, or a 4th category: ‘angel’, meaning they had achieved best practice. The 35 design-phase participants preferred 3 levels of opportunity, associating them with a ‘good, better, best’ paradigm, and appreciated the self-efficacy boost of knowing they were already achieving best practice.
Four demarcations were also used to categorise ML opportunities using feature importance, measured through average gain across all splits where a feature was used. The higher a feature’s importance, the higher an individual’s sensitivity to that feature. Feature-importance thresholds were set so each individual had at least 4 opportunities in the last 2 columns of the LoF matrix, equivalent to 2- and 3-star rules-based opportunities. The ML model was more discriminating, e.g. double the number of 3-star opportunities in Figure 5, and 9 ‘angels’, compared to 1 ‘angel’ in the rules-based model.
Participants were informed that the vertical axis was the expert conclusion based on their data and objective, and the horizontal axis was the ML assessment of their individual sensitivity using the same inputs. The supporting rationale, combining the individual’s validated data with sleep-research insights, was displayed in the user interface, using the rules-based model logic.
Participants were instructed to select interventions that they: understood; were able to control; and were motivated to change [43]. They were required to give themselves an ‘adequate sleep opportunity’ for 7 hours of sleep, based on SE (of time asleep while in bed). They were given decisional freedom to choose up to 3 other interventions on the LoF matrix. Once they chose their interventions, they were asked to set specific targets (e.g. get up between 6:45 and 7:15 every day, including weekends).
Figure 6: Example tracker daily text and dashboard.
The user-interface tracker provided daily feedback over the course of 1 month of intervention on how participants were doing against these actions. The user-interface tracker presented a simple dashboard accompanied by an automated text message each day. The user could investigate each metric, using the same data visualisation as they had used to validate their data (now colour-coded relative to their target) and longer-term views of the changes.
3.3 Measuring the ML leap of faith
Intervention choice, compliance, and sleep outcomes were recorded, enabling three relative-trust metrics to be assessed: DIRTI, DAFTI, and DOTI. We illustrate each using pilot data; see Figure 7 for how these metrics relate to the types of trust discussed in the literature.
Figure 7: DIRTI, DAFTI and DOTI target demonstrated trust (intention-setting; action-follow); and deserved trust. Icons by Angriawan Ditya Zulkamain, Arif Hariyanto, Karen Tyler and PEBIAN from thenounproject.com CC BY 3.0.
Demonstrated Intention Relative Trust Index (DIRTI)measures how users demonstrate trust in ML through their intentions. Using two study participants on the left-hand side of Figure 8 as an example, D chose 1 action with an ML score of above zero on their LoF horizontal axis; all other choices only scored highly on the rules axis. D’s average rules score was 2 (pale blue) and ML 0.5 (dark blue). Most of B’s choices scored high on the ML axis; some scored lower on rules, so they averaged 2.5 on ML and 2.25 on rules. B placed more relative trust in ML than D: B’s DIRTI, calculated by dividing their ML score by their rules score, is 1.1; D’s 0.25, as shown on the right-hand side of Figure 8. Although all participants could have chosen actions with at least equal ML and rules scores, most relied more on rules than on ML when they set intentions; only 2 had DIRTI >1.
Figure 8: DIRTI: average ML intervention priority (dark blue) divided by average rule priority (light blue) yields the relative trust demonstrated in ML at intention on right. Blue lines are averages.
Figure 9: B’s mean light exposure from wake (CT0) to bedtime during baseline (left-hand side) and intervention (right-hand side), in blue (shaded area is 95th percentile). Red line is period mean.
Figure 10: Indexed percentage improvement vs baseline in blue; deterioration vs baseline in red.
Demonstrating trust requires action, not just intention, which is what Demonstrated Action Follow-through Index (DAFTI)measures. Some interventions required major changes to participants’ daily routines. One commented ‘I don’t need a sleep coach; I need a life coach’ when planning how to follow through on their intentions. For example, as illustrated in Figure 9: B made a big increase in their daily exposure in intervention compared to baseline, in red, and shifted a lot of it into the afternoon. Figure 10 shows participant follow-through, comparing the last 7 days of intervention to baseline. The last 7 days was used because most participants had to work towards target behaviour in stages over the 4 weeks. The proliferation of blue indicates most showed some follow-through. D was very successful on caffeine; B on natural light. Some, in red, failed to follow through: D on 2; B’s and E’s bedroom temperatures increased, but they made other changes to cool their bodies down.
The DAFTI metric reflects action follow-through by discarding an intervention score if a participant fails to improve on baseline. For example, D acted on their caffeine intention, but failed to implement a relaxation routine, their only intervention with a positive ML score. As shown in Figure 11, D’s average ML score therefore dropped to 0, and their rules score dropped to 1.1 (light blue) because they also failed to follow through on their natural-light intention. Their relative demonstrated trust in ML (DAFTI), dividing their ML score by their rules score, was 0. B followed through on all but one action, and their relative trust increased to 1.2. Three participants now showed more trust in ML than rules (DAFTI >1): F did a great job on evening light, a high-scoring ML intervention, and not so well on re-calibrating natural light levels, a high-scoring rules intervention.
Figure 11: DAFTI: average ML priority after follow-through (dark blue) divided by average rule priority (light blue) yields the relative trust demonstrated in ML after follow-through (mid-blue). If the participants showed any improvement vs baseline, the score was included. Blue lines are averages.
Our ML models were designed to be trustworthy, so the outstanding deserved-trust question is whether there is a relationship between the models’ prioritisation of recommendations and the outcomes delivered, i.e. did participants who implemented high-scoring interventions get better outcomes? Was the risk worth it? The Deserving Of Trust index (DOTI)measures whether the trust was worth it, reflecting the constraints faced once the participant acted. There will be a placebo effect at play, which can be considerable, but there is no reason to assume it would not align to demonstrated trust.
Demonstrated trust is shown on an absolute basis on the left of Figure 12 – i.e. the correlation (${r}^{2}$) between each individual’s average ML or rules score and their percentage increase in TST and SWS. We compare baseline to the last seven days of intervention because most participants worked towards target behaviour in stages over four weeks of intervention. In this pilot, the strongest relationship was for ML-interventions at follow-through (dark blue): TST ${r}^{2}$ increases from 0.58 at intention to 0.76; SWS from 0.22 to 0.54. An increase is consistent with a trustworthy model because recommendations only work if people act on them. Relative DOTI, the ratio of the absolute ML score and the rules score, is shown on the right. Trust was more deserved in the ML algorithm for TST than for SWS, which may have been influenced by lower SWS model accuracy. At follow-through, our ML model deserved more trust than the rules-based one; had only intention been considered, the SWS rules-based model would have (incorrectly) been seen to deserve more trust.
Figure 12: Absolute DOTI on the left: TST at top and SWS on bottom: correlation (${r}^{2}$) between intention and follow-through scores, and % improvement in last 7 days’ sleep end points relative to baseline. Relative DOTI is on the right, calculated by dividing the ML metric by the rules-based one.
4 Discussion
Contribution
The proposed architecture allows the ML leap of faith to be discerned and measured, and therefore to be managed over time. Experimental design lessons include:
More human agency and control is afforded by data validation and objective-function specification, enabled by separating data and objective function from model reasoning and data visualisation. Users build intrinsic trust by pre-validating data in an accessible format, and deciding the model objective.
An explicit reference standard, created through a collectively-developed rules-based model, enables experts to test it against their mental model a priori. They should do so until they intrinsically trust it, so it becomes a verified point of reference. Industry or regulatory bodies could maintain the standard as critical infrastructure – e.g. for assurance audits or to train up new experts. Maintenance should include regularly stress testing it to keep assumptions and use cases up-to-date, and verify ongoing trust. Many domains already use rules-based models, which should evolve into a reference standard.
A LoF matrix visually and accessibly contrasts the output of the reference model with the ML model’s a posteriori output, using the same data and objective function, thus bringing the ML leap of faith into sharp focus. This provides real transparency to model users: they should be inclined to trust where models agree, and understand a leap of faith is only required where they disagree. Experts can stress test the areas of disagreement, updating the rules-based model with new insights or to capture drift, thus narrowing the leap of faith as an aspect of operations management. Different ML algorithms will require different leaps of faith, because a priori levels of expert insight differ, as do model accuracies.
By being explicit about the ML leap of faith, we can assess demonstrated and deserved trust. For the first time, our metrics measure trust in terms of actions and outcomes. If trust is deserved, so relative DOTI increases over time; DIRTI and DAFTI should also increase. These simple metrics allow trust comparisons across models, and flag the need for investigation if they trend down. Industry regulators could track companies’ ML-model-trust metrics, especially DOTI, to understand trends. Continuous audit processes using these metrics would preserve ML algorithm IP whilst building public assurance.
Limitations
The matrix and metrics may be perceived as insufficiently technical. They are accessible by design, so they can be used and understood by experts, businesses, and regulators as well as computer/ data scientists. They need to be applied at scale to reach empirical conclusions about trust for different users/ domains/ situations, and what drives it, objectively. The proposed metrics do not measure attitudinal trust or model trustworthiness. Instead they measure how attitude is reflected in decisions, actions, and outcomes, which drive ML adoption. Because our proposed metrics only measure intentions, actions and outcomes, they could be applied to an untrustworthy system and used to encourage adoption. If GDPR (or similar) is not followed, trust metrics could be inappropriately sold and/ or used to discriminate against people based on their general or specific trust inclination.
Participant time for data validation can be substantial if data does not readily align with perception. Our approach may be most suitable for professionals who use model recommendations often, so their time commitment pays off as they focus on where there is a leap of faith to be overcome. The time required to develop a reference standard from scratch can be considerable, requiring: substantial programmer subject-matter education; individuals who can translate expert knowledge into rules; and time from experts on each iteration, and to test the resulting model for different use cases. However, experts may resist allocating this time to ML, especially if they have a high opinion of their own capability and/or are concerned about their roles being replaced or diminished. The data-validation approach is not universally accessible: our participants had at least one university degree and understood complex, novel concepts and images. Work is needed to evaluate accessibility for other levels of educational attainment. Prioritisation, unlike direct regression/ classification comparisons, requires calibration if it is to be included in a LoF matrix: different approaches should be tested.
Conclusion
The proposed architecture is a more practical, rigorous way of building trust in an ML model than trying to explain it to people who have a different (probably unarticulated) mental model. Critical novel architecture features include: user data validation and objective-function specification; an explicit reference standard; a LoF matrix; and demonstrated- and deserved-trust metrics.